OpenStudy (anonymous):

This is my equation and I am having trouble figuring it out... 10(x+y) to the 0 power.. can someone help?

Parth (parthkohli):

Okay, so do you know what anything to the 0 power is?

OpenStudy (anonymous):

yes it is 1

OpenStudy (anonymous):

anything to the power of zero is 1, unless the base is zero. so when ever x+y not equal to zero. it is 1

OpenStudy (anonymous):

okay so from there my equation would be 10(1)?

OpenStudy (anonymous):

depends if you want \[10(x+y)^{0}\] it is 10(1), or if you whole power(including 10 also) it is 1

OpenStudy (anonymous):

okay my equation is 10(x+y)^{0}

OpenStudy (anonymous):

by the way you have written the equation, it seems 10 is excluded from the zero power.so it is 10..

OpenStudy (anonymous):

so i would not do this : 10(x+y)^{0} 10(x ^{0}+y ^{0}) which would be: 10(1+1) and then: 10(2) and then 20

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

hey sorry, that is wrong

OpenStudy (anonymous):

my answer?

OpenStudy (anonymous):

see you can not take the power inside like that, it should be \[(x+y)^{0}\]

OpenStudy (anonymous):

which is 1, so 10*1 is 10

OpenStudy (anonymous):

if the question is \[(10(x+y))^{0}\] then it is 1

OpenStudy (anonymous):

okay I understand now, shoot you wanna come be my teacher cause my teacher isnt any good.. lol do you have time to help me with few more?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

sweet! okay here we go!

OpenStudy (anonymous):

simplify. write the answer using positive exponents only.

OpenStudy (anonymous):

\[(8x ^{-4}y ^{-2}z ^{3}/2xy ^{-2}z ^{-3})^{-1}\]

OpenStudy (anonymous):

ok first think \[(x ^{a})^{b} =x ^{ab}\]

OpenStudy (anonymous):

exponents are multiplied here, correct

OpenStudy (anonymous):

\[(xy)^{a} = x ^{a}*y ^{a}\]

OpenStudy (anonymous):

yah perfect, multiply the exponents

OpenStudy (anonymous):

okay give me a second im gonna post up my work.

OpenStudy (anonymous):

Okay so if we break everything up into peices.. (excuse the spelling im not an english major I PROMISE :-) )

OpenStudy (anonymous):

lol.. neither me

OpenStudy (anonymous):

we have: \[8^{1*-1}=-8\] \[x ^{-4*-1}=x ^{4}\] \[y ^{-2*-1}=y ^{2}\] \[z ^{3*-1} = z ^{-3}\] divided by: \[2^{1*-1}=-2\] \[x ^{1*-1}=x ^{-1}\] \[y ^{-2*-1}=y ^{2}\] \[z ^{-3*-1}=z ^{3}\] so in the end we have \[-8x ^{4}y ^{2}z ^{-3}\div-2x ^{-1}y ^{2}z ^{3}\]

OpenStudy (anonymous):

\[8^{-1}= \frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

and from there we move them accordingly such as.. idk if i am right i am just throwing out what i think...

OpenStudy (anonymous):

o okay now im lost haha

OpenStudy (anonymous):

okay i will tell you step wise..

OpenStudy (anonymous):

nvm i see where u got that

OpenStudy (anonymous):

everything you have done is correct except -2 and -8, they are 1/8 and 1/2

OpenStudy (anonymous):

okay now so to eliminate the negative exponents dont I just subtract then flip them except for the y's because they equal each other out right.. because \[y ^{2}-y ^{2}=0\] correct?

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{a} } = x ^{-a}\]

OpenStudy (anonymous):

correct.. so in the end my equation minus the 8&2 should be this: \[x ^{5}\div z ^{6}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[\frac{ x ^{4} }{ x ^{-1} } = x ^{4-(-1)}\]

OpenStudy (anonymous):

= \[x ^{5}\]

OpenStudy (anonymous):

awesome.. thats it

OpenStudy (anonymous):

what about 8 &2

OpenStudy (anonymous):

do they flip spots as well to make them even

OpenStudy (anonymous):

just 8*x^5/2*z^6 =4*x^5/z^6

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

8*?

OpenStudy (anonymous):

2*?

OpenStudy (anonymous):

sorry my bad, x^5/4*z^6

OpenStudy (anonymous):

because 2/8 = 1/4

OpenStudy (anonymous):

\[\frac{ x ^{5} }{ 4*z ^{6} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay not to be nosey or anything but what is ur name?

OpenStudy (anonymous):

charan, and you??

OpenStudy (anonymous):

paige.. r u a b or g

OpenStudy (anonymous):

hehe.. b

OpenStudy (anonymous):

how old im a g

OpenStudy (anonymous):

20 , which class r u in??

OpenStudy (anonymous):

im not trying to freak you out sorry.. i am 21 almost 22 and i swear if you are younger than me i feel like a retard

OpenStudy (anonymous):

i go to lone star college so my math is called 0308 its like 1 or 2 levels down from college math but not the very first one.. if that makes any sense

OpenStudy (anonymous):

hehe.. nothing there to feel anything. i am sure you are good at something else, and yah im nt freaked out :-)

OpenStudy (anonymous):

lol... i m usually pretty good at math... i just missed like a whole week of classes.. so I am so far behind and I have a test when I get back and I am already failing this class cause my teacher is a feather.. and I have to PASS!!! if I dont they take my financial aid and i will lose my extra money for my house and I have a 6 mnth old daughter I cant afford to lose my house.. but anyways.. so can u still help me with some more since you are super smart.. and by the way your user name is a total lie.. your not a loser your freaking smart!

OpenStudy (anonymous):

hehe.. im not super smart. yah i can help you.

OpenStudy (anonymous):

you rock!! and well your smarter than me that is for sure... i wish this board thing worked better...

OpenStudy (anonymous):

okay i am doing something relly quick so if i lose you please come find me again cause I DESPERATLY need help lol

OpenStudy (anonymous):

okay sorry im back i had to fix my wireless

OpenStudy (anonymous):

whats your next question?

OpenStudy (anonymous):

okay hold on let me get it.. I am going through this review trying to figure out which ones i know and which ones I dont.. where r u from.. if you dont mind me asking while i find the next problem

OpenStudy (anonymous):

im from tomball/cypress area i dont know if you know where that is but it is in texas

OpenStudy (anonymous):

okay next problem...

OpenStudy (anonymous):

ok, im from india.

OpenStudy (anonymous):

ooo okyy.. wow so far away.. what time is it there?

OpenStudy (anonymous):

9pm

OpenStudy (anonymous):

wait hold on we did this on i forgot to correct it hold please i do have more problems tho

OpenStudy (anonymous):

okay got it number 5... Directions: Simplify and write positive exponents only. the problem is: \[7^{-1}+3^{-1}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 7 } +\frac{ 1 }{ 3 } = \] \[\frac{ 10^{1} }{ 21^{1} }\]

OpenStudy (anonymous):

okay so its like the other problem we did

OpenStudy (anonymous):

yah.. i dont think any more simplification can be done

OpenStudy (anonymous):

no it cant..

OpenStudy (anonymous):

okay you ready for another one.. these next like 2 or 3 im gonna give you the problem and the answer and can u tell me if they r right please?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay the first one goes like this: directions: Add Problem: \[(3x ^{2}y-4xy+7)+(-2x ^{2}y+12xy-9)\] Answer: \[x ^{2}y+8xy-2\]

OpenStudy (anonymous):

correct answer

OpenStudy (anonymous):

Sweet.. Okay next.. Unless you have to go.. i dont want to keep you up.. cause its like 11 am here

OpenStudy (anonymous):

i wnt sleep this early, you can ask your questions

OpenStudy (anonymous):

Okay, your a life saver, just let me know if you need to go!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay here is another for you... directions: multiply question: \[(x-6)(-5x-9)\] Answer: \[-5x ^{2}+21x+54\] Question: \[(-3x ^{2}+8y)(-3x ^{2}-7y)\] Answer: \[9x ^{4}-3x ^{2}y-56y ^{2}\]

OpenStudy (anonymous):

perfect.. both are correct :)

OpenStudy (anonymous):

Oky as you are grading me I am working the problems.. I am trying to go as fast as i can.. but okay I am stuck again.. Directions: Multiply Question: (x+1/9)(x-1/9) Answer: I am Lost.. I got \[x ^{2}-1/81\]

OpenStudy (anonymous):

thats it :)

OpenStudy (anonymous):

but the -1/81 the doesnt make sense i guess to me

OpenStudy (anonymous):

no.. why do you feel so, (a+b)*(a-b) = \[a ^{2}-b ^{2}\]

OpenStudy (anonymous):

okay.. i trust you i guess it just looks weird to me but your the math genius so i shall not argue with u

OpenStudy (anonymous):

this is too much, that equation i studied in my 6th class :-) syllabus in India and US are right

OpenStudy (anonymous):

Okay I am pretty sure this one is right but I just want to double check... Directions: Multiply Question: \[(5xy ^{2}+6x)(5xy ^{2}-6x)\] Answer: \[25x ^{2}y ^{4}-36x ^{2}\]

OpenStudy (anonymous):

right :-)

OpenStudy (anonymous):

you studied this in 6TH GRADE!!!?????

OpenStudy (anonymous):

yah 6th or 7th grade :)

OpenStudy (anonymous):

holy pellet.. i feel really dumb now.. why do yall want to come over here yll education system is much better than ours it seems.. yall r way more advanced... i dont mean that in a racial way at all i swear..